Hi, I am having trouble with part b of this question.
A piece of wire is 60cm long and is bent to form a right-angled triangle having the hypotenuse of 25cm with the other sides being x and y.
a) Express y in terms of x.
y=60-25-x
y=35-x
b) Use the Pythagoras Theorum to write an equation in x. Hence find the length of the other two sides of the triangle.

Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

A piece of wire is 60cm long and is bent to form a right-angled triangle having the hypotenuse of 25cm with the other sides being x and y.
a) Express y in terms of x.
y=60-25-x
y=35-x
b) Use the Pythagoras Theorum to write an equation in x. Hence find the length of the other two sides of the triangle.
------
x^2 + (35-x)^2 = 25^2
x^2 + 1225 - 70x + x^2 = 625
2x^2 - 70x + 600 = 0
x^2 - 35x + 300 = 0
(x-20)(x-15) = 0
---
x = 20 cm then y = 15 cm
OR
x = 15 cm then y = 20 cm
----------
Cheers,
Stan H.
-----------------

Answer by Alan3354(69443)

A piece of wire is 60cm long and is bent to form a right-angled triangle having the hypotenuse of 25cm with the other sides being x and y.
a) Express y in terms of x.
y=60-25-x
y=35-x
b) Use the Pythagoras Theorum to write an equation in x. Hence find the length of the other two sides of the triangle.
------
x^2 + y^2 = 25^2
x + y = 35
y = 35 - x
---
x^2 + (35-x)^2 = 625
2x^2 - 70x + 600 = 0
x^2 - 35x + 300 = 0
(x-15)*(x-20) = 0
x = 15, 20
y = 20, 15